OpenStudy (anonymous):

The function f(t) = 3 cos(pi over 6t) + 5 represents the tide in Blastic Sea. It has a maximum of 8 feet when time (t) is 0 and a minimum of 2 feet. The sea repeats this cycle every 12 hours. After nine hours, how high is the tide? 12 feet 5 feet 4.5 feet 2.5 feet

OpenStudy (amistre64):

let t=9 ....

OpenStudy (amistre64):

and i believe its not 'over 6t' but 'over 6' then times t

OpenStudy (anonymous):

here ill right out what it looks like before i start to work it out

OpenStudy (anonymous):

\[f(t) = 3\cos(\frac{ \pi }{ 6 })+5\]

OpenStudy (anonymous):

supposed tobe a t in the parenthesis

OpenStudy (amistre64):

\[f(t) = 3\cos(\frac{ \pi }{ 6 }t)+5\]or \[f(t) = 3\cos(\frac{ \pi }{ 6t})+5\]

OpenStudy (anonymous):

first one

OpenStudy (amistre64):

at any rate, let t=9 and the solution just solves itself

OpenStudy (anonymous):

k one sec

OpenStudy (anonymous):

i got 5 feet

OpenStudy (amistre64):

5 it is then

OpenStudy (anonymous):

thank you

OpenStudy (amistre64):

yw